OpenStudy (anonymous):

An equation is shown below: 5(2x – 3) = 5 Part A: How many solutions does this equation have? (4 points) Part B: What are the solutions to this equation? Show your work. (6 points)

pooja195 (pooja195):

Distribute

pooja195 (pooja195):

5 x 2x=? 5 x -3 =?

OpenStudy (anonymous):

7 ,-8

pooja195 (pooja195):

You are supposed to multiply

OpenStudy (anonymous):

10 , -15

pooja195 (pooja195):

5(2x−3)=5 (5)(2x)+(5)(−3)=5(Distribute) 10x+−15=5 10x−15=5 Step 2: Add 15 to both sides. 10x−15+15=5+15 10x=20 Step 3: Divide both sides by 10.

pooja195 (pooja195):

so whats 20/ 10 ?

OpenStudy (anonymous):

10 @pooja195

pooja195 (pooja195):

No.

pooja195 (pooja195):

20 divided by 10 = ?